OpenStudy (anonymous):

Let \(r\) and \(s\) be two real or complex numbers and m and n be two integers \ such that \[r + s = m \\ r s = n\] and let \(k\ge 1\) be an integer.Show that \[r^k + s^k\] is an integer. You can prove it by the methods used to prove http : // openstudy.com/study #/updates/5498 ce78e4b0b8a54fae51fb

OpenStudy (solomonzelman):

strange having to prove this. idk how to prove it. you have a sum of numbers and both numbers are integers raised to the power of integers, so why wouldn't they be integers?

OpenStudy (solomonzelman):

I know this is not enough, but I was just wondering.

OpenStudy (blacksteel):

r and s are not necessarily integers.

OpenStudy (solomonzelman):

yes but you can substitute m and n for them, and that will make it then.

OpenStudy (solomonzelman):

I wish I can help, sorry.

OpenStudy (blacksteel):

Let's start by defining:\[r^k + s^k = (r+s)^k - \sum_{i=1}^{k-1}\left(\begin{matrix}k \\ i\end{matrix}\right)r^is^{k-i}\]Since we know k is an integer greater than 0, this is just a simple reexpression of Pascal's Rule. By the given parameters of the problem, r+s is an integer, so (r+s)^k is an integer, so we simply have to prove that the value of the summation is an integer. If k is even, then when i = k/2, i=k-i. Since the value of rs is an integer, the value of \[\left(\begin{matrix}k/2 \\ k/2\end{matrix}\right)r^{k/2}*s^{k/2}\] will be an integer. For each other value j of i such that j < k/2, there will be a related value k-j of i. Then consider the arbitrary case of one such pair of values: \[\left(\begin{matrix}k \\ j\end{matrix}\right)r^{j}*s^{k-j} + \left(\begin{matrix}k \\ k-j\end{matrix}\right)r^{k-j}*s^{j} =\]\[\left(\begin{matrix}k \\ j\end{matrix}\right)r^j*s^j*s^{k-2j} + \left(\begin{matrix}k \\ j\end{matrix}\right)r^j*s^j*r^{k-2j} =\]\[(\left(\begin{matrix}k \\ j\end{matrix}\right)r^j*s^j)(s^{k-2j} + r^{k-2j})\] Since rs is an integer and r+s is an integer, this pair of terms has an integer value. Since each term of the summation is either a member of such a pair, or has an integer value, the value of the summation is an integer and as such the value of \[r^k + s^k\]is an integer.

OpenStudy (blacksteel):

Oops, I left out a step. Since j is an integer and j < k/2, k - 2j is an integer greater than 0. Thus, we can use induction to prove that \[s^{k-2j}+r^{k-2j}\] being a lower-exponent case of the same problem, is an integer. Since s^1 + r^1 = s + r is an integer, we have our base case.

OpenStudy (anonymous):

If you look at my post @Blacksteel in http://openstudy.com/users/eliassaab#/updates/5498ce78e4b0b8a54fae51fb you can write an easier proof.

OpenStudy (blacksteel):

I'm not clear on why the assumption that r and s are roots of a polynomial equation holds. The original problem you based your solution on explicitly states that, but using a trivial case like r = 2, s = 3, m = 5, n = 6, all of the assumptions of this problem hold and r and s are clearly not roots of \(x^2 + x - 1 = 0\)

OpenStudy (anonymous):

No, they are the solutions of a similar but not the same quadratic. But, you are right in the fact that I should mention r and s could be real or complex number. I have edited the problem to reflect this.

OpenStudy (blacksteel):

That wasn't my point at all - your solution assumes that r and s are the roots of the SAME polynomial. If they aren't, you can't construct an equation with \(r^{k-2}\) and \(s^{k-2}\) having the same coefficient, which means you can't use induction. I haven't done due diligence to state that your solution isn't applicable with modifications here, but it definitely doesn't translate directly.

OpenStudy (anonymous):

Take the quadratic \[ x^2 - m x + n =0\] This equation has two real or complex solutions r and s, such that r+s=m and r s =n

OpenStudy (anonymous):

Do you know the sum and product formula for the two roots of a quadratic?

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

The sum (S) and the product (P) of real or complex roots of \[ a x^2 + b x + c=0\\ \text { are }\\ S= -\frac b a\\ P = \frac c a \]

OpenStudy (blacksteel):

Okay, that's the part I was missing. (I know the sum and product formulas and how to construct a polynomial with given roots, but you didn't state in the original proof that that was what you used and I wasn't sure how you got there, and therefore how to generalize it).